# Dual thermal pseudo-critical features in a spin-1/2 Ising chain with twin-diamond geometry

### v1

@@ III.3 Pseudo-critical temperature
$T_{p_{1}}\approx\frac{3J_{0}+2J_{1}-J_{2}+2(h_{1}-h_{0})}{2k_{\mathrm{B}}\ln(2)}.$ (32)
Similarly one can obtain the second pseudo-critical temperature between $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$, which becomes
@@ IV.1 Entropy and specific heat
Panels (c) and (d) present the corresponding specific heat $C(T)$, which develops two continuous maxima at $T_{p_{1}}$ and $T_{p_{2}}$. The logarithmic insets make the locations of these maxima particularly visible. The first maximum marks the crossover between the ordered FI region and the partially frustrated $\mathrm{FR}_{1}$ configurations, while the second reflects the competition between $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$. We can observe that the peak at $T_{p_{1}}$ is higher and sharper than the second peak at $T_{p_{2}}$. This behavior is consistent with the pseudo-critical features reported in the literature[4, 6, 13, 12].
@@ V Conclusion
In this work we have presented an exact analysis of the coupled twin-diamond chain (CTDC), a decorated one-dimensional Ising system inspired by the magnetic structure of $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$. The model contains two inequivalent sublattices, allowing for distinct local fields, and includes three exchange couplings that control the interplay between nodal spins and internal dimers. Using the decoration-iteration mapping[29, 30, 31] and the resulting $4\times 4$ transfer matrix, we obtained analytic expressions for the eigenvalues governing the full thermodynamics of the system.

### v2

@@ III.3.3 Pseudo-critical temperature
$T_{p_{1}}\approx\frac{3J_{0}+2J_{1}-J_{2}+2(h_{1}-h_{0})}{2k_{\mathrm{B}}\ln(2)},$ (36)
which accurately reproduces the low-temperature behavior of $T_{p_{1}}$ obtained in Eq. (33). In a similarly manner, the pseudo-critical temperature associated with the $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$ competition is estimated as
@@ IV.1 Entropy and specific heat
Panels (c) and (d) present the corresponding specific heat $C(T)$, which develops two continuous maxima at $T_{p_{1}}$ and $T_{p_{2}}$. The logarithmic insets make the locations of these maxima particularly visible. The first maximum marks the crossover between the ordered FI region and the partially frustrated $\mathrm{FR}_{1}$ configurations, while the second reflects the competition between $\mathrm{FR}_{1}$ and $\mathrm{FR}_{2}$. We can observe that the peak at $T_{p_{1}}$ is higher and sharper than the second peak at $T_{p_{2}}$. This behavior is consistent with the pseudo-critical features reported in the literature(sergio18, ; Rojas2019, ; katarina, ; verissimo, ).
@@ IV.2 Magnetic properties
IV.3 Eigenvalues properties and correlation length
Figure 7(a) shows the temperature dependence of the relative eigenvalues $|\lambda_{j}|/\lambda_{1}$ for the parameter set $J_{2}/J_{1}=1.24$, $J_{0}/J_{1}=-0.01$, $\mu_{\mathrm{B}}g_{0}B/J_{1}=1.1$, and $g_{1}/g_{0}=0.7$. Although the eigenvalues satisfy the ordering $\lambda_{1}>\lambda_{2}>\lambda_{3}>\lambda_{4}$, the subleading eigenvalue in modulus is not $\lambda_{2}$ but rather $|\lambda_{4}|$ (see appendix A). This behavior is illustrated by the red curve $|\lambda_{4}|/\lambda_{1}$, while the green and blue curves correspond to $\lambda_{2}/\lambda_{1}$ and $\lambda_{3}/\lambda_{1}$, respectively. The ratio $|\lambda_{4}|/\lambda_{1}$ exhibits a broad plateau between the two pseudo-critical temperatures, indicating that the same low-energy sector remains dominant throughout this interval. In contrast, the ratio $\lambda_{2}/\lambda_{1}$ displays two pronounced extrema located precisely at the pseudo-critical temperatures $T_{p_{1}}$ and $T_{p_{2}}$, reflecting the successive competition between distinct thermodynamic sectors.
For the same parameters, Fig. 7(b) shows the quantities $\xi_{j}=\bigl[\ln(\lambda_{1}/|\lambda_{j}|)\bigr]^{-1}$ as functions of temperature on a logarithmic scale. Since $|\lambda_{4}|$ is the largest subleading eigenvalue in modulus, the spectral correlation length is given by $\xi=\xi_{4}=\bigl[\ln(\lambda_{1}/|\lambda_{4}|)\bigr]^{-1}$. However, this quantity does not exhibit signatures of the pseudo-critical behavior. By contrast, the quantity $\xi_{2}=\bigl[\ln(\lambda_{1}/\lambda_{2})\bigr]^{-1}$, although not a correlation length in the strict spectral sense, clearly signals both pseudo-critical temperatures $T_{p_{1}}$ and $T_{p_{2}}$. In general, the spectral correlation length is determined by
$\xi=\left[\ln\left(\frac{\lambda_{1}}{\max\{|\lambda_{2}|,|\lambda_{3}|,|\lambda_{4}|\}}\right)\right]^{-1},$ (38)
whereas pseudo-critical behavior is associated with near-degeneracies between the leading eigenvalues that govern the dominant thermodynamic sectors.
Figure 7: (a) Temperature dependence of the relative eigenvalues $|\lambda_{j}|/\lambda_{1}$ for the fixed parameters $J_{2}/J_{1}=2.48$, $J_{0}/J_{1}=-0.01$, $\mu_{{\rm B}}g_{0}B/J_{1}=1$, and $g_{1}/g_{0}=1.3$. (b) Temperature dependence of the quantities $\xi_{j}=\ln(\lambda_{1}/\lambda_{j})^{-1}$ for the same set of parameters as in panel (a).
Although universality and quasicritical exponents have been identified for a broad class of one-dimensional decorated models (Rojas2019, ), a quantitative scaling analysis of correlation lengths is not pursued here. In contrast, the quantity $\xi_{2}=[\ln(\lambda_{1}/\lambda_{2})]^{-1}$ exhibits anomalous critical behavior, but a detailed analysis of its associated exponents lies beyond the scope of the present work.
The emergence of two pseudo-critical temperatures in the present model is closely related to the coupling of two diamond units within the same chain. However, increasing the number of diamond units or coupling additional chains does not, in general, guarantee the appearance of further pseudo-critical temperatures. In the coupled twin-diamond chain, the internal structure of the unit cell supports two frustrated manifolds, $\mathrm{FR_{1}}$ and $\mathrm{FR}_{2}$, with different extensive degeneracies, leading to successive entropy-driven free-energy crossings as temperature increases and hence to two well-separated pseudo-transitions. Additional pseudo-transitions may arise only when the enlarged unit cell stabilizes new competing low-energy sectors with distinct residual entropies that survive the coupling, consistently with the criterion discussed in Ref. (bjp20, ).
@@ V Conclusion
In this work we have presented an exact analysis of the coupled twin-diamond chain (CTDC), a decorated one-dimensional Ising system inspired by the magnetic structure of $\mathrm{Cu}_{2}(\mathrm{TeO}_{3})_{2}\mathrm{Br}_{2}$. The model contains two inequivalent sublattices, allowing for distinct local fields, and includes three exchange couplings that control the interplay between nodal spins and internal dimers. Using the star-triangle mapping(dec-trns, ; ddec-trns, ; jozef-dctrns, ) and the resulting $4\times 4$ transfer matrix, we obtained analytic expressions for the eigenvalues governing the full thermodynamics of the system.
@@
Appendix A Pairwise structure of the eigenvalues
Here we clarify the algebraic structure of the four eigenvalues of the transfer matrix and show that two eigenvalues associated with the Perron branch, namely $\lambda_{1}$ and $\lambda_{2}$ given in Eqs. (15) and (16), are necessarily real, while the remaining eigenvalues may be either real or form a complex-conjugate pair.
A.1 Pairwise algebraic structure of the spectrum
The exact diagonalization yields the four eigenvalues in the form
$\lambda_{1}$ $=c+\frac{R_{1}+\sqrt{R_{-}}}{2},\qquad\lambda_{2}=c+\frac{R_{1}-\sqrt{R_{-}}}{2},$ (39)
$\lambda_{3}$ $=c-\frac{R_{1}-\sqrt{R_{+}}}{2},\qquad\lambda_{4}=c-\frac{R_{1}+\sqrt{R_{+}}}{2},$ (40)
with
$c=\frac{\mathfrak{r}+\mathfrak{t}}{4},\qquad R_{1}=\sqrt{2y-a_{2}},$ (41)
and
$R_{\mp}=-2y-a_{2}\mp\frac{2a_{1}}{R_{1}}.$ (42)
Here all coefficients are real, and the auxiliary variable $y$ is chosen as a real root of the resolvent cubic.
Equation (39) shows that the spectrum naturally decomposes into two algebraic pairs: $(\lambda_{1},\lambda_{2})$, controlled by the radicand $R_{-}$, and $(\lambda_{3},\lambda_{4})$, controlled by $R_{+}$. Any complex eigenvalues can therefore arise only within one of these pairs.
A.2 Reality of $\lambda_{1}$ and $\lambda_{2}$
The transfer matrix has strictly positive entries and is primitive. By the Perron-Frobenius theorem (cuesta, ), the dominant eigenvalue $\lambda_{1}$ is real, strictly positive, and nondegenerate, and is given by the first expression in Eq. (39). If $R_{-}<0$, then $\sqrt{R_{-}}$ would be purely imaginary, causing $\lambda_{1}$ to acquire a nonzero imaginary part, which is impossible. Therefore $R_{-}\geqslant 0$, and both $\lambda_{1}$ and $\lambda_{2}$ are real on the thermodynamic branch.
A.3 Remaining eigenvalues and possible complex conjugate pair
No analogous restriction applies to the second radicand $R_{+}$. If $R_{+}<0$, then $\sqrt{R_{+}}$ is purely imaginary, and Eq. (39) immediately yields $\lambda_{4}=\lambda_{3}^{*}$, so that $\lambda_{3}$ and $\lambda_{4}$ form a complex-conjugate pair. If $R_{+}\geqslant 0$, all four eigenvalues are real.
Thus, on the thermodynamic branch, the spectrum always contains at least two real eigenvalues, $\lambda_{1}$ and $\lambda_{2}$, while the remaining pair may be real or complex conjugate, depending on the Hamiltonian parameters.
A.4 Magnitude of eigenvalues
Although the eigenvalues $\lambda_{j}$ are labeled according to their algebraic expressions, their ordering on the real axis is parameter dependent. Even when all eigenvalues are real, the second-largest eigenvalue in magnitude need not coincide with $\lambda_{2}$; depending on parameters, it may instead correspond to $\lambda_{3}$ or $\lambda_{4}$. Consequently, subleading behavior is governed by the eigenvalue with the largest modulus among $\lambda_{j\neq 1}$, rather than by a fixed algebraic branch.